integrating factor by inspection (2x^2 + y)dx + (x^2y - x)dy = 0


1
Expert's answer
2022-05-09T18:21:47-0400

integrating factor by inspection:

"\left(2x^2+y\right)dx+\left(x^2y-x\right)dy=0"

Solution:

"\left(2x^2dx+x^2ydy\right)+\left(ydx-xdy\right)=0",

"\frac{x^2\left(2dx+ydy\right)}{x^2}-\frac{\left(xdy-ydx\right)}{x^2}=0" ,

"\left(2dx+ydy\right)-\frac{\left(xdy-ydx\right)}{x^2}=0" ,

"d\left(2x+\frac{y^2}{2}\right)-d\left(\frac{y}{x}\right)=0" ,

"d\left(2x+\frac{y^2}{2}-\frac{y}{x}\right)=0" ,

"2x+\frac{y^2}{2}-\frac{y}{x}=C" ,

Answer: "U\left(x,y\right)=2x+\frac{y^2}{2}-\frac{y}{x}=C" , integrating factor is "\mu\left(x\right)=\frac{1}{x^2}"